OpenStudy (anonymous):

(5x)(3x)(6x2 – 1) 90x3 – 15x 48x3 – 8x 48x4 – 8x2 90x4 – 15x2

OpenStudy (anonymous):

it is solving for polynomials

jhonyy9 (jhonyy9):

hope so much that you can multiplie sure right ,yes ?

jhonyy9 (jhonyy9):

so i think the last one will be right

jhonyy9 (jhonyy9):

because 5*3*6=90 and x*x=x^2 *x^2 =x^4 and you will continue it and will get the right answer sure right good luck bye